OpenStudy (anonymous):

dy/dx of y=(sqrt(x)-1)/(sqrt(x)+1)

OpenStudy (anonymous):

\[y=(\sqrt{x-1})/(\sqrt{x+1)}\]

OpenStudy (anonymous):

is it?

OpenStudy (anonymous):

@mathhelpkl

OpenStudy (anonymous):

it's \[y=\sqrt{x}-1/\sqrt{x}+1\]

OpenStudy (anonymous):

ok..ok

OpenStudy (anonymous):

Do you knw the answer?

OpenStudy (anonymous):

not quite...I'm really close though I think

OpenStudy (anonymous):

where are you now on your solution?

OpenStudy (anonymous):

I've gotten to \[y \prime=1/(\sqrt{x}+1) \] but I don't think that's quite right

OpenStudy (anonymous):

* the bottom is squared

OpenStudy (anonymous):

i am getting \[y'=1/(\sqrt{x}*(\sqrt{x+1})^2)\]

OpenStudy (anonymous):

how did you get \[\sqrt{x}\] on the bottom

OpenStudy (anonymous):